OpenStudy (anonymous):

x=4cos^2t y=2sin2t Solve please ?

OpenStudy (hba):

Cartiesian co-ordinate sys ?

OpenStudy (anonymous):

you just need to find the circle or something

OpenStudy (anonymous):

when I solve it is (x^2)/4 +(y^2)/4=cos^2t,and that is not goog because I do not know is it the pharabola or circle or somthingelse I'm from Serbia,because of that my english is bad

OpenStudy (anonymous):

I just have the cordinate,evrything else is fixed

OpenStudy (anonymous):

I solve it by myself

OpenStudy (anonymous):

It's an ellipse actually. Since the coefficients in front of the 2 aren't the same. If you had: \[x=\alpha \cos(\beta t); y=\gamma \sin(\beta t)\] Then for: \[\alpha=\gamma\] You have a circle and it is traced out with: \[0 \le t \le \frac{2 \pi}{\beta}\] If you have something with: \[\alpha \ne \gamma\] Then you have an ellipse because when cos(beta t) is maximum (one) you are at ALPHA. But when sin(beta t) is maximum (one) you are at GAMMA. So you have something like: \[\frac{x}{\alpha}=\cos(\beta t); \frac{y}{\gamma}=\sin(\beta t); \cos^2(\beta t)+\sin^2(\beta t)=\left( \frac{x}{\alpha} \right)^2+\left( \frac{y}{\gamma} \right)^2=1\] Which is the equation for an ellipse with major/minor axes alpha and gamma.